A line which contains these points: (-3,3) &
(-8,-2) would have a slope of _________?
(simplify your answer)

Found 2 solutions by DJ_Sammie, checkley71:
Answer by DJ_Sammie(24)

m=3-(-2)/-3-(-8)
m=3+2/-3+8
m=5/5
m=1
the slope is 1

Answer by checkley71(8403)

SLOPE = (Y2-Y1)/(X2-X1) THUS
(-2-3)/(-8+3)=-5/-5=1 THE SLOPE IS 1